40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 2 00:27:20 2019, gautam, Update, ALS, Green ALS resurrection ALSnoiseComparison.pdfALS_TR_RIN.pdfGreemAM.pdf
    Reply  Mon Dec 2 18:20:29 2019, gautam, Update, ALS, EY uPDH post mixer LPF EY_uPDH_LPF.pdfEY_uPDH_OLTF.pdf
       Reply  Mon Dec 2 20:18:18 2019, rana, Update, ALS, EY uPDH post mixer LPF 
          Reply  Tue Dec 3 18:15:42 2019, gautam, Update, ALS, EY uPDH post mixer LPF 
             Reply  Tue Dec 3 21:28:24 2019, gautam, Update, ALS, EY uPDH post mixer LPF uPDH.pdf
       Reply  Tue Dec 3 00:10:15 2019, Koji, Update, ALS, EY uPDH post mixer LPF 
          Reply  Tue Dec 3 00:51:25 2019, gautam, Update, ALS, EY uPDH post mixer LPF 
Message ID: 15058     Entry time: Mon Dec 2 00:27:20 2019     Reply to this: 15059
Author: gautam 
Type: Update 
Category: ALS 
Subject: Green ALS resurrection 

Attachment #1 - comparison of phase tracker servo angle fluctuations for the green beat vs IR beat.

  • To convert to Hz, I used the PT servo calibration detailed here.
  • This is only a function of the delay line length and not the signal strength, so shouldn't be affected by the difference in signal strength between the IR and green beats.
  • For the green beat - I divided the measured spectra by 2 to convert the green beat frequency fluctuations into equivalent IR frequency fluctuations.
  • There is no whitening before digitization. I believe the measured spectra are dominated by ADC noise above ~50 Hz. See this elog for the frequency discriminant as a funtion of signal strength, so 5uV/rtHz ADC noise would be ~2 Hz/rtHz for a -5dBm signal, which is what I expect for the Y beat, and ~0.5 Hz/rtHz for a +5dBm signal, which is what I expect for the X beat. Hence the brown (Green beat, XARM) being lower than the green trace (IR beat, XARM) isn't real, it is just because of my division of 2. So I guess that calibration factor I applied is misleading.
  • I did not yet check the noise in the other configuration - arm lengths controlled using ALS, and POX/POY as the OOL sensors. To be tried tonight.

Attachment #2 - RIN of the DCPDs.

  • I noticed that over 10s of seconds, the GTRY level was fluctuating by ~5%. 
  • This was much more than any drift seen in the GTRX level.
  • Measuring the RIN on the DCPDs (Thorlabs PDA36A) supports this observation (spectra were divided by DC value to convert into RIN units).
  • There is ~120uW (1.6 VDC, compatible with 30dB gain setting) incident on the GTRX PD, and ~6uW (170 mVDC, compatible with 40dB gain setting) incident on the GTRY PD.
  • Not sure what is driving this drift - I don't see any coherence with the IR TRY signal, so doesn't seem like it's the cavity.

Characterization of the green beat setup [past numbers]:

  • With some patient alignment effort (usual near-field/far-field matching), I was able to recover the green beat signals.
  • Overall, the numbers I measured today are consistent with what was seen in the past when we had the ability to lock using green ALS.
  • The mode-matching between the PSL and AUX green beams are still pretty abysmal, ~40-50%. The mode shapes are clearly different, but for now, I don't worry about this.
  • I saw some strong AM of the beat signal (for both EX and EY beats) while I was looking at it on a scope, see Attachment #3. This AM is not visible in the IR beat, not sure what to make of it. The frequency of the AM is ~1 MHz, but it's hard to nail this down because the scope doesn't have a very long buffer, and I didn't look at the frequency content on the Agilent (yet).

o BBPD DC output (mV), all measured with Fluke DMM

             XARM   YARM 
 V_DARK:     +1.0    +2.0
 V_PSL:      +8.0    +13.0
 V_ARM:      +157.0  +8.0


o BBPD DC photocurrent (uA)
I_DC = V_DC / R_DC ... R_DC: DC transimpedance (2kOhm)
 I_PSL:       3.5    5.5
 I_ARM:      78.0    3.0


o Expected beat note amplitude
I_beat_full = I1 + I2 + 2 sqrt(e I1 I2) cos(w t) ... e: mode overlap (in power)
I_beat_RF = 2 sqrt(e I1 I2)

V_RF = 2 R sqrt(e I1 I2) ... R: RF transimpedance (2kOhm)

P_RF = V_RF^2/2/50 [Watt]
     = 10 log10(V_RF^2/2/50*1000) [dBm]

     = 10 log10(e I1 I2) + 82.0412 [dBm]
     = 10 log10(e) +10 log10(I1 I2) + 82.0412 [dBm]

for e=1, the expected RF power at the PDs [dBm]
 P_RF:      -13.6  -25.8


o Measured beat note power (measured with oscilloscope, 50 ohm input impedance)      
 P_RF:      -17.95dBm (80 mVpp)  -28.4dBm (24mVpp)   (40MHz and 42MHz)  
    e:        37%                    55  [%]                                             

I also measured the various green powers with the Ophir power meter (filter off): 

o Green light power (uW) [measured just before PD, does not consider reflection off the PD]
 P_PSL:       18    24
 P_ARM:       400     13

The IR beat is not being made at the moment because I blocked the PSL beam entering the fiber.

Attachment 1: ALSnoiseComparison.pdf  84 kB  | Hide | Hide all
ALSnoiseComparison.pdf
Attachment 2: ALS_TR_RIN.pdf  44 kB  | Hide | Hide all
ALS_TR_RIN.pdf
Attachment 3: GreemAM.pdf  194 kB  Uploaded Mon Dec 2 11:57:10 2019  | Hide | Hide all
GreemAM.pdf
ELOG V3.1.3-